OpenStudy (anonymous):

Rewrite in simplest radical form 1/x^−3/6. Show each step of your process.

OpenStudy (anonymous):

i know how to get to -1/2 when you simplify the exponent, but i dont know where to go from there

jhonyy9 (jhonyy9):

so for example sqrtx = x^1/2 where 2 is the index of radical - this will help you ?

jhonyy9 (jhonyy9):

courage - this is easy

OpenStudy (anonymous):

im confused

jhonyy9 (jhonyy9):

so you know again that 1/x = x^(-1) yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

-3/6 simplifie by 3 and will get -1/2 yes ?

OpenStudy (anonymous):

yes i have that

jhonyy9 (jhonyy9):

so than there are 1/x^(-1/2) = x^(-(-1/2)) so what will result now ?

OpenStudy (anonymous):

would it be x^1/2?

jhonyy9 (jhonyy9):

yes what mean sqrtx yes ?

jhonyy9 (jhonyy9):

do you understand it now ?

OpenStudy (anonymous):

kinda, how do i solve it in simplest radical form though?

jhonyy9 (jhonyy9):

simplest radical form will be sqrtx

OpenStudy (anonymous):

oh, ok thank you

jhonyy9 (jhonyy9):

np was my pleasure - anytime - good luck bye